# A distribution-guided Mapper algorithm

### v2

@@ Abstract
Background: The Mapper algorithm is an essential tool to explore the shape of data in topology data analysis. With a dataset as an input, the Mapper algorithm outputs a graph representing the topological features of the whole dataset. This graph is often regarded as an approximation of a Reeb graph of data. The classic Mapper algorithm uses fixed interval lengths and overlapping ratios, which might fail to reveal subtle features of data, especially when the underlying structure is complex.
Results: In this work, we introduce a distribution-guided Mapper algorithm named D-Mapper, that utilizes the property of the probability model and data intrinsic characteristics to generate density-guided covers and provides enhanced topological features. Moreover, we introduce a metric accounting for both the quality of overlap clustering and extended persistent homology to measure the performance of Mapper-type algorithms. Our numerical experiments indicate that the D-Mapper outperforms the classical Mapper algorithm in various scenarios. We also apply the D-Mapper to a SARS-COV-2 coronavirus RNA sequence dataset to explore the topological structure of different virus variants. The results indicate that the D-Mapper algorithm can reveal both the vertical and horizontal evolution processes of the viruses. Our code is available at https://github.com/ShufeiGe/D-Mapper.
@@ 1 Introduction
In recent decades, machine learning methods have become popular tools to discover valuable information from data. These methods typically focus on prediction tasks by establishing a mapping between responses and predictors. The shape of data may reveal important information in data from a new perspective and provide alternative insights to supplement prediction tasks, which is often overlooked in the learning of mapping. To be specific, the shape of data refers to the manifold formed by the support set of data distribution. The shape of data reveals the data distribution, reflecting the spatial correlations and dependency structures among data points. Such information is crucial in clustering and feature extraction tasks. However, many existing methods cannot fully utilize such information.
@@
Topology data analysis (TDA) is a specific area of investigating data shapes based on topological theory. Topology is a powerful tool to study the shape of objects by finding invariants that remain unchanged under continuous deformations. These invariants can reflect objects’ intrinsic features. TDA attempts to identify and utilize these topology invariants in data analysis. The Mapper algorithm is a simple but powerful TDA tool for visualizing and clustering data. The algorithm constructs a graph in which each vertex represents a cluster, and each edge represents the two adjacent clusters that share some elements. The colour and size of vertices are defined by users and can provide more information. For example, the size of vertices and the colour depth can represent the average value and the number of elements in the cluster. The Mapper algorithm was first proposed in [1] and has been applied in various domains, such as social media [2], fraud detection [3], and evolutionary computation [4]. The Mapper algorithm is especially suitable for biology data, which are usually complex and high-dimensional [5]. For example, the Mapper algorithm was applied to breast cancer transcriptional microarray data and successfully identified two subgroups with $100\%$ survival rate [6]. It is also used to analyze transcriptional programs that control cellular lineage commitment and differentiation during development, and the proposed scTDA identified four transient states over time [7]. Exploring the conformation space of proteins is another important task in computational biology, and the Mapper algorithm has been successfully applied to analyze intermediate conformations, the results are closely consistent with experimental findings [8].
@@
The classic Mapper algorithm requires a filter function $f:X\rightarrow\mathbb{R}^{p}$ to project data $X$ onto the Euclidean space $\mathbb{R}^{p}$ and a set of open cover $\mathcal{U}$ on the projection data.In most of the literature, $p$ is set to $1$, i.e. the filter function $f$ maps data from a high dimensional space to a real line. Different filter functions and open covers may result in different outputs, so it is necessary to select them carefully. Improper filter functions or open covers may not accurately reveal data shapes, resulting in poor overlap clustering. A filter function is chosen primarily based on features of interest and also depends on specific applications. Determining the optimal parameters of the model, such as the overlapping rates and interval lengths, usually requires extensive manual tuning and experiments. As pointed out in [9], cover choices lead to a broad range of possible Mapper outputs, resulting in many different graph depictions and clustering results. Unreasonable choices may cause the disappearance of certain topological features.
@@
In this work, alternatively, we propose a distribution-guided Mapper algorithm to relax the restriction of regular covers and overlapping ratios. Our proposed algorithm utilizes the property of the probability model and data intrinsic characteristics to generate distribution-guided covers and provides enhanced topological features. Unlike the classic Mapper algorithm, our proposed D-Mapper does not rely on predefined static overlapping rates and interval lengths, but instead, we fit projected data to a mixture distribution model and generate flexible covers automatically. Moreover, the performance of the Mapper algorithm is hard to evaluate due to the complexity of the output graph structure. Methods in most of the literature only evaluate it from the clustering perspective and overlook the quality of topological structure [11, 12]. In this paper, we introduce a metric that can quantitatively and objectively evaluate the performance of the Mapper algorithm from both clustering and topological aspects. We use the extended persistent homology as a tool to capture topological signatures, and then we distinguish noise and signal on the diagram by constructing a confidence set on the diagram, the signal rate can be used as a metric to evaluate the quality of topological structure. The simulation studies indicate that the D-Mapper algorithm outperforms the classic Mapper algorithm. The remainder of the paper is organized as follows. We briefly review the classic Mapper algorithm and introduce our D-Mapper and evaluation metric in the section 2. Section 3 focuses on the comparison of the classic Mapper and the D-Mapper on simulated datasets and the application of the proposed D-Mapper algorithm on a real-world SARS-COV-2 RNA sequence dataset. We discuss and summarize our work in sections 4 and 5.
@@ 2.1 Basic notions
The union of the simplices of $K$ constitutes the underlying space of $K$, denoted as $|K|$, which inherits from the topology space of $\mathbb{R}^{p}$. Thus the geometric simplicial complex can also be regarded as a topological space.
@@
Suppose $\mathcal{U}=(u_{i}),i\in I$ is a collection of open subset of a topological space $X$, then we say $\mathcal{U}$ is an open cover of $X$ if $X=\bigcup_{i\in I}u_{i}$.
@@
Let $\mathcal{U}=(u_{i}),i\in I$ be an open cover of a paracompact space $X$ by open sets such that the intersection of any sub-collection of the $u_{i}$’s is either empty or contractible. Then, $X$ and the nerve C($\mathcal{U}$) are homotopy equivalent.
@@ 2.2 Mapper algorithm
The Mapper algorithm is an important tool to construct the discrete version of Reeb graphs which encode connected information of the support manifold [15]. Algorithm 1 depicts the classic Mapper algorithm. Firstly, the original data is projected onto a real line by a user-specified filter function (Algorithm 1, line 1). To construct a cover $\mathcal{U}$ on the projected data, the number of components $n$ and overlapping ratio $p$ should be chosen carefully (Algorithm 1, line 2). The improper choice of these parameters may lead to failure estimation of the shape of data. With these two parameters, a cover with $n$ regular intervals can be obtained (Algorithm 1, line 3). Then, the inverse image of the cover $\mathcal{U}$ is achieved on the original data, generating some hypercubes in the original data space. This process is often called pulling back (Algorithm 1, line 4). Finally, clustering data points within each hypercube. Each cluster corresponds to a vertex in the Mapper graph. If two vertices share any elements, an edge is added between these two vertices (Algorithm 1, line 5).
@@
3:Construct a cover $\mathcal{U}=(u_{i}),i=1...n$ on projected data $f(X)$ based on the parameter $n$ and $p$.
4:Pull back the intervals of projected data to the original space, $f^{-1}(\mathcal{U})$.
@@
The classic Mapper algorithm is simple but powerful for exploring and visualizing data, while the selection of parameters $n,~{}p$ involves extensive manual tuning and the algorithm is sensitive to these parameters. In addition, the Mapper algorithm’s flexibility is often restricted by regular intervals and fixed overlapping ratios, which may hinder the discovery of complex data structures.
@@ 2.3 D-Mapper
The restriction of regular intervals with fixed overlapping ratios is one of the major limitations of the Mapper algorithm. We propose a distribution-guided Mapper algorithm to generate flexible covers to reflect the underlying data structures. Our proposed method automatically chooses the overlapping ratios based on the distribution of the projection data and produces more flexible covers to reveal the data shapes more accurately. The key idea of our algorithm is to fit the projected data with a mixture probability model. Each component in the mixture model can be viewed as an interval, and the probability (likelihood) of each data point assigned to each interval can be explicitly calculated. Once we get the mixture distribution of the projected data, we can create intervals based on the distribution in many ways.
Here we introduce a simply way to construct intervals based on the quantile $\alpha$ of the probability distribution. With proper selection, the $\alpha$ quantile interval of each component of the mixture model can automatically produce some overlaps. Figure 1 shows the idea of naturally producing intervals by quantile $\alpha$. This attribute provides a natural scheme for constructing flexible covers on the projected data. The specific procedures are as follows:
1) Choose an appropriate number of intervals $n$ and a quantile $\alpha$. The number of components naturally matches the number of intervals.
@@
3) Intervals are determined by the $\alpha$ quantile intervals of each component of the mixture model.
Figure 1: An illustration of intervals produced by the D-Mapper algorithm. The deep blue line represents the probability density function of each component in the GMM. The shallow blue dashed line presents the probability density function of the GMM. The orange lines are intervals that produced naturally by symmetry quantile $\alpha$. (a) When $\alpha=0.01$, there are overlaps between adjacent intervals. (b) When $\alpha=0.1$, there is a gap between the first and second interval. $\alpha$ controls the overlap of intervals, it should be chosen carefully.
@@
By incorporating a mixture model into the Mapper algorithm, it can produce flexible intervals and overlaps. We call this distribution-guided Mapper algorithm as D-Mapper algorithm. With these flexible intervals, we can construct the output graph similar to the classic Mapper algorithm: pulling back these intervals to the original data space, then clustering data points within each interval separately, all sub-groups constitute the vertices of the nerve, and adding an edge between two vertices if there are any shared data points. Algorithm 2 gives details of the D-Mapper.
Algorithm 2 D-Mapper
@@
2:Choose a component number $n$ and quantile $\alpha$.
@@
9:The collection of intervals $\mathcal{U}=(u_{i}),i=1...n,$ is a cover on projected data $f(X)$.
@@
11:Cluster on the refined cover and build the nerve by clustering result.
The quantile $\alpha$ controls the overlapping ratios: larger $\alpha$ leads to lower overlapping ratios. Distributions of components in mixture models are often heterogeneous, resulting in flexible quantile intervals automatically. Proper $\alpha$ selection is crucial to ensure all points are covered. The Mapper algorithms usually require pairwise overlap (i.e. each interval has overlap with its neighbours)[15, 10, 11, 12]. For D-Mapper, the pairwise overlap property can also be preserved, as shown in Figure 1 (a). However, larger $\alpha$ may result in disjoint intervals, Figure 1 (b) gives an example of disjoint intervals caused by improper large $\alpha$. Thus, in this work, we propose a method to find the upper bound of $\alpha$ that guarantees the pairwise overlap property. By establishing this upper bound, we also significantly narrow the potential range of $\alpha$ values, thereby facilitating parameter tuning. Denote $F_{i}^{-1}$ the inverse of the cumulative density function of the $i$th ordered component, thus the $i$th interval $u_{i}$ is given by $[F_{i}^{-1}(\alpha/2),F_{i}^{-1}(1-\alpha/2)]$, $i=1,\ldots,n$. The idea of this method is to find the upper bound $\alpha^{\prime}$, that guarantees the intersection of each paired adjacent intervals is not empty, $u_{i}\cap u_{i+1}\neq\emptyset$, i.e., $F_{i}^{-1}(1-\alpha/2)\geq F_{i+1}^{-1}(\alpha/2)$, $i=1,2,\ldots,n-1$. Notice that although pairwise overlap is a good property for a Mapper graph, this property is not necessary for all situations. One can construct a good Mapper graph with some paired adjacent intervals being non-overlap as long as the union of intervals can cover all data points. As shown in Figure 6 (a) in Section 3, since there are no points between the two disjoint circles, disjoint paired adjacent intervals should be presented in the corresponding Mapper output graph. We add one threshold parameter $\alpha^{*}$ to allow for disjoint paired adjacent intervals (lines 6-10 of Algorithm 3). In practice, we suggest to set $\alpha^{*}=0.005$. Algorithm 3 describes how to find the upper bound of $\alpha$.
@@
10:$\alpha^{\prime}=min\{\alpha_{i}\}_{i\in\boldsymbol{J}}$.
@@ 2.4 Evaluation metric
The silhouette coefficient ($SC$) is often used as a measure to evaluate the quality of (overlap) clustering [17]. The $SC$ assesses how well the clusters are separated, how close the clusters are, and is stable for overlap clustering. Suppose we have $n$ samples in dataset $D$ that could be divided into $k$ clusters: $C_{1},...,C_{k}$. For any two data points $x,~{}x^{\prime}\in D$, the compactness of the cluster to which $x$ belongs can be defined as:
@@
where $d(x,x^{\prime})$ represents the distance between $x$ and $x^{\prime}$, $|C_{i}|$ is the number of data points in cluster $i$. The degree of separation between $x$ and other clusters can be computed as:
$b(x)=\mathop{\min}_{C_{j}:j\neq i}\bigg{\{}\frac{\sum_{x^{\prime}\in C_{j}}d(x,x^{\prime})}{|C_{j}|}\bigg{\}},$ (2)
And the $SC$ of $x$ is:
$SC(x)=\frac{b(x)-a(x)}{\max\{a(x),b(x)\}},$ (3)
@@
The $SC$ reflects the clustering quality of data points, but it does not evaluate the topology structure of a Mapper output graph. Figure 2 shows an example of different Mapper graphs on the same dataset. The left panel has a higher $SC$ but a poor topological feature. Conversely, the right panel has a lower $SC$ but a good topological structure. Figure 2 (a) shows that small overlap $p$ makes one of the circles disconnected, however, more separable clusters lead to a high $SC$. Therefore, a metric that evaluates both clustering and topological structure is needed. Although, the topological information encoded in Mapper has been well studied theoretically [15]. There is no practical metric to evaluate the topology structure of Mapper graphs. Most existing work evaluates the topology structure of Mapper graphs based on the clustering results only. The extended persistence diagram has been proven to be a powerful tool for capturing topological signatures of Mapper graph [18]. In this manuscript, we alleviate it to provide a simple quantitative metric for evaluating the topology structure of Mapper output.
@@
Figure 2: An example of the classic Mapper algorithm on the same dataset but outputs different graphs. The dataset is shown in Figure 6 (a), the classical Mapper is implemented, and the clustering algorithm is DBSCAN with a radius of $0.5$ and a minimum of samples $3$. (a) The output graph when $n=12,p=0.01$ and its $SC=0.283$, $SC_{adj}=0.521$. This graph has a higher $SC$ but a poor topological structure. (b) The output graph when $n=12,p=0.1$ and its $SC=0.246$, $SC_{adj}=0.812$. This graph has a lower $SC$ but a good topological structure.
@@
The extended persistence diagram is an extension of the persistence diagram, the extended persistence diagram uses both sub-level sets and super-level sets as a filtration, therefore it contains more information than the original persistence diagram [19]. In practical, given a graph with a function defined on its nodes, we can compute the extended persistence diagram that reflects the topological features of this graph. For a Mapper graph, the function on a node can be naturally defined as the mean value of points in the node. The points on the diagram can be regarded as signatures of the Mapper graph. However, these signatures are not always meaningful, points near the diagonal are often seen as noise. To distinguish the real signal from the noise, one simple yet effective method is to calculate the confidence interval using bottleneck bootstrap and then identify the noise based on this interval[9]. Bottleneck bootstrap is an effective way to compute confidence set on a persistence diagram [20]. It uses bootstrap samples to get a bootstrap Mapper graph and then calculates the bottleneck distance from the original Mapper graph. Repeat this step many times, and an approximate distribution of bottleneck distances can be obtained, then we can easily calculate the confidence set with this approximate distribution. More details of the bottleneck bootstrap algorithm are shown in Appendix B.
@@
We introduce a metric to evaluate the quality of the extended persistence diagram depending on the confidence set, we call this metric the topological signal rate, denoted as $TSR$. The $TSR$ is defined as the number of real signal points divided by the total number of points on the extended persistence diagram, serving as a quantitative indicator of the quality of the extend persistence diagram.
@@
An example of an extended persistence diagram and its $TSR$ is illustrated in Figure 3. The gray area is the confidence set estimated by the bottleneck bootstrap. Points inside are noises and those outside are signal points. The $TSR$ in this example is $0.25$, indicating poor quality of the corresponding Mapper graph. To get a proper evaluation of the Mapper graph, we can combine the $TSR$ with $SC$ through weighted averaging. We call this metric an adjusted silhouette coefficient.
Figure 3: An example of the extended persistence diagram. There are a total of $8$ points in the diagram, the gray area is computed by the bottleneck bootstrap, and points inside this area are noise. Thus, $6$ points are noises and $2$ points are signals, the $TSR$ is $0.25$.
@@
Figure 2 compares the value of $SC_{adj}$ with the original $SC$. The original $SC$ overlooks the topological structure of Mapper graph, while the $SC_{adj}$ can give a reasonable evaluation. In the next section, we show more examples to validate the effectiveness of $SC_{adj}$.

### v3

@@ Abstract
Background: The Mapper algorithm is an essential tool for exploring the data shape in topological data analysis. With a dataset as an input, the Mapper algorithm outputs a graph representing the topological features of the whole dataset. This graph is often regarded as an approximation of a Reeb graph of a dataset. The classic Mapper algorithm uses fixed interval lengths and overlapping ratios, which might fail to reveal subtle features of a dataset, especially when the underlying structure is complex.
Results: In this work, we introduce a distribution-guided Mapper algorithm named D-Mapper, which utilizes the property of the probability model and data intrinsic characteristics to generate density-guided covers and provide enhanced topological features. Moreover, we introduce a metric accounting for both the quality of overlap clustering and extended persistent homology to measure the performance of Mapper-type algorithms. Our numerical experiments indicate that the D-Mapper outperforms the classic Mapper algorithm in various scenarios. We also apply the D-Mapper to a SARS-COV-2 coronavirus RNA sequence dataset to explore the topological structure of different virus variants. The results indicate that the D-Mapper algorithm can reveal both the vertical and horizontal evolutionary processes of the viruses. Our code is available at https://github.com/ShufeiGe/D-Mapper.
@@ 1 Introduction
In recent decades, machine learning methods have become popular tools to discover valuable information from data. These methods typically focus on prediction tasks by establishing a mapping between responses and predictors. The data shape may reveal important information in data from a new perspective and provide alternative insights to supplement prediction tasks, which is often overlooked in the learning of mapping. To be specific, the data shape refers to the manifold formed by the support set of data distribution. The data shape reveals the data distribution, reflecting the spatial correlations and dependency structures among data points. Such information is crucial in clustering and feature extraction tasks. However, many existing methods cannot fully utilize such information.
@@
Topology data analysis (TDA) is a specific area of investigating data shapes based on topological theory. Topology is a powerful tool for studying the shape of objects by finding invariants that remain unchanged under continuous deformations. These invariants can reflect objects’ intrinsic features. TDA attempts to identify and utilize these topology invariants in data analysis. The Mapper algorithm is a simple but powerful TDA tool for visualizing and clustering data. The algorithm constructs a graph in which each vertex represents a cluster, and each edge represents the two adjacent clusters that share some elements. The colour and size of vertices are defined by the users and can provide additional information. For example, the size of vertices and the colour depth can represent the average value and the number of elements in the cluster. The Mapper algorithm was first proposed in [1] and has been applied in various domains, such as social media [2], fraud detection [3], and evolutionary computation [4]. The Mapper algorithm is especially suitable for biology data, which are usually complex and high-dimensional [5]. For example, the Mapper algorithm was applied to breast cancer transcriptional microarray data and successfully identified two subgroups with $100\%$ survival rate [6]. It is also used to analyze transcriptional programs that control cellular lineage commitment and differentiation during development, and the proposed scTDA identified four transient states over time [7]. Exploring the conformation space of proteins is another important task in computational biology, and the Mapper algorithm has been successfully applied to analyze intermediate conformations, the results are closely consistent with experimental findings [8].
@@
The classic Mapper algorithm requires a filter function $f:X\rightarrow\mathbb{R}^{p}$ to project data $X$ onto the Euclidean space $\mathbb{R}^{p}$ and a set of open cover $\mathcal{U}$ on the projected data. In most of the literature, $p$ is set to $1$, i.e. the filter function $f$ maps data from a high dimensional space to a real line. Different filter functions and open covers may result in different outputs, so it is necessary to select them carefully. Improper filter functions or open covers may not accurately reveal data shapes, resulting in poor overlap clustering. A filter function is chosen primarily based on the features of interest and also depends on specific applications. Determining the optimal parameters of the model, such as the overlapping rates and interval lengths, usually requires extensive manual tuning and experiments. As pointed out in [9], cover choices lead to a broad range of possible Mapper outputs, resulting in many different graph depictions and clustering results. Unreasonable choices may cause the disappearance of certain topological features.
@@
In this work, alternatively, we propose a distribution-guided Mapper (D-Mapper) algorithm to relax the restriction of regular covers and overlapping ratios. Our proposed algorithm utilizes the property of the probability model and data intrinsic characteristics to generate distribution-guided covers and provides enhanced topological features. Unlike the classic Mapper algorithm, our proposed D-Mapper does not rely on predefined static overlapping rates and interval lengths. Instead, we fit the projected data to a mixture distribution model and automatically generate flexible covers. Moreover, evaluating the performance of the Mapper algorithm is challenging due to the complexity of the output graph structure. Most methods in literature only evaluate it from a clustering perspective, overlooking the quality of topological structure [11, 12]. In this work, we introduce a metric that can quantitatively and objectively evaluate the performance of the Mapper algorithm from both clustering and topological aspects. We use the extended persistent homology as a tool to capture topological signatures, and distinguish noise from signal on the diagram by constructing a confidence set on the diagram. The signal rate can be used as a metric to evaluate the quality of topological structure. The simulation studies indicate that the D-Mapper algorithm outperforms the classic Mapper algorithm. The remainder of the paper is organized as follows. We briefly review the classic Mapper algorithm and introduce our D-Mapper and evaluation metric in the section 2. Section 3 focuses on comparing the classic Mapper with the D-Mapper on simulated datasets and applying the proposed D-Mapper algorithm to a real-world SARS-COV-2 RNA sequence dataset. We discuss and summarize our work in Sections 4 and 5.
@@ 2.1 Basic notions
The union of the simplices of $K$ constitutes the underlying space of $K$, denoted as $|K|$, which inherits from the topology space of $\mathbb{R}^{p}$. Thus, the geometric simplicial complex can also be regarded as a topological space.
@@
Suppose $\mathcal{U}=(u_{i}),i\in I$ is a collection of open subset of a topological space $X$, then we say $\mathcal{U}$ is an open cover of $X$ if black $X=\bigcup_{i\in I}u_{i}$.
@@
Let $\mathcal{U}=(u_{i}),~{}i\in I$ be an open cover of a paracompact space $X$ by open sets such that the intersection of any sub-collection of the $u_{i}$’s is either empty or contractible. Then, $X$ and the nerve C($\mathcal{U}$) are homotopy equivalent.
@@ 2.2 Mapper algorithm
The Mapper algorithm is an important tool to construct the discrete version of Reeb graphs, which encode connected information of the support manifold [15]. Algorithm 1 depicts the classic Mapper algorithm. Firstly, the original data is projected onto a real line by a user-specified filter function (Algorithm 1, line 1). To construct a cover $\mathcal{U}$ on the projected data, the number of components $n$ and overlapping ratio $p$ should be chosen carefully (Algorithm 1, line 2). The improper choice of these parameters may lead to failure estimation of the data shape. With these two parameters, a cover with $n$ regular intervals can be obtained (Algorithm 1, line 3). Then, the inverse image of the cover $\mathcal{U}$ is achieved on the original data, $f^{-1}(\mathcal{U})$, generating some hypercubes in the original data space. This process is often called pulling back (Algorithm 1, line 4). Finally, clustering data points within each hypercube. Each cluster corresponds to a vertex in the Mapper graph. If two vertices share any elements, an edge is added between these two vertices (Algorithm 1, line 5).
@@
3:Construct a cover $\mathcal{U}=(u_{i}),i=1...n$ on the projected data $f(X)$ based on the parameters $n$ and $p$.
4:Pull back the intervals of the projected data to the original space, $f^{-1}(\mathcal{U})$.
@@
The classic Mapper algorithm is simple but powerful for exploring and visualizing data, while the selection of parameters $n,~{}p$ involves extensive manual tuning, and the algorithm is sensitive to these parameters. In addition, the Mapper algorithm’s flexibility is often restricted by regular intervals and fixed overlapping ratios, which may hinder the discovery of complex data structures.
@@ 2.3 D-Mapper
The restriction of regular intervals with fixed overlapping ratios is one of the major limitations of the Mapper algorithm. We propose a distribution-guided Mapper algorithm to generate flexible covers to reflect the underlying data structures. Our proposed method automatically chooses the overlapping ratios based on the distribution of the projected data and produces more flexible covers to reveal the data shapes more accurately. The key idea of our algorithm is to fit the projected data with a mixture probability model. Each component in the mixture model can be viewed as an interval, and the probability (likelihood) of each data point assigned to each interval can be explicitly calculated. Once we get the mixture distribution of the projected data, we can create intervals based on the distribution in many ways.
Here, we introduce a simple way to construct intervals based on a confidence level of $1-\alpha$ for a given probability distribution. With a proper selection, $1-\alpha$ confidence intervals of each component of the mixture model can automatically produce some overlaps. Figure 1 shows the idea of naturally producing intervals given a confidence level of $1-\alpha$. This attribute provides a natural scheme for constructing flexible covers on the projected data. The specific procedures are as follows.
1) Choose an appropriate number of intervals $n$ and a confidence level of $1-\alpha$. The number of components naturally matches the number of intervals.
@@
3) Intervals are determined by $1-\alpha$ confidence intervals of each component of the mixture model.
Figure 1: An illustration of intervals produced by the D-Mapper algorithm. The deep blue line represents the probability density function of each component in the GMM. The shallow blue dashed line presents the probability density function of the GMM. The orange lines are intervals that are produced naturally given a confidence level of $1-\alpha$. (a) When $\alpha=0.01$, there are overlaps between adjacent intervals. (b) When $\alpha=0.1$, there is a gap between the first and second intervals. $\alpha$ controls the overlap of intervals, and it should be chosen carefully.
@@
By incorporating a mixture model into the Mapper algorithm, it can produce flexible intervals and overlaps. We call this distribution-guided Mapper algorithm the D-Mapper algorithm. With these flexible intervals, we can construct the output graph similar to the classic Mapper algorithm: pulling back these intervals to the original data space, then clustering data points within each interval separately, all sub-groups constitute the vertices of the nerve, and adding an edge between two vertices if there are any shared data points. Algorithm 2 gives details of the D-Mapper.
Algorithm 2 The D-Mapper aglorithm
@@
2:Choose a component number $n$ and a confidence level $1-\alpha$.
@@
9:The collection of intervals $\mathcal{U}=(u_{i}),i=1...n,$ is a cover on the projected data $f(X)$.
@@
11:Cluster on the refined cover and build the nerve with the clustering result.
The confidence level $1-\alpha$ controls the overlapping ratios: a larger $\alpha$ value leads to lower overlapping ratios. The distributions of components in mixture models are often heterogeneous, automatically resulting in flexible intervals. Proper $\alpha$ selections are crucial to ensure all points are covered. The Mapper algorithms usually require pairwise overlapping (i.e. each interval overlaps with its neighbours)[15, 10, 11, 12]. The D-Mapper algorithm can also reserve the pairwise overlap property, as shown in Figure 1 (a). However, a larger $\alpha$ value may result in disjoint intervals, Figure 1 (b) gives an example of disjoint intervals caused by an improper large $\alpha$ value. Thus, in this work, we propose a method to find the upper bound of $\alpha$ that guarantees the pairwise overlap property. By establishing this upper bound, we also significantly narrow the potential range of $\alpha$ values, thereby facilitating parameter tuning. Denote $F_{i}^{-1}$ the inverse of the cumulative density function of the $i$th ordered component, thus the $i$th interval $u_{i}$ is given by $[F_{i}^{-1}(\alpha/2),F_{i}^{-1}(1-\alpha/2)]$, $i=1,\ldots,n$. The idea of this method is to find the upper bound $\alpha^{\prime}$, that guarantees the intersection of each paired adjacent intervals is not empty, $u_{i}\cap u_{i+1}\neq\emptyset$, i.e., $F_{i}^{-1}(1-\alpha/2)\geq F_{i+1}^{-1}(\alpha/2)$, $i=1,2,\ldots,n-1$. Notice that although pairwise overlapping is a good property for a Mapper graph, this property is not necessary for all situations. One can construct a good Mapper graph with some paired adjacent intervals being non-overlap as long as the union of intervals can cover all data points. As shown in Figure 6 (a) in Section 3, since there are no points between the two disjoint circles, disjoint paired adjacent intervals should be presented in the corresponding Mapper output graph. We add one threshold parameter $\alpha^{*}$ to allow for disjoint paired adjacent intervals (lines 6-10 of Algorithm 3). In practice, we suggest to set $\alpha^{*}=0.005$. Algorithm 3 describes how to find the upper bound of $\alpha$.
@@
10:$\alpha^{\prime}=min\{\alpha_{i}^{\prime}\}_{i\in\boldsymbol{J}}$.
@@ 2.4 Evaluation metric
The silhouette coefficient ($SC$) is often used as a measure to evaluate the quality of (overlap) clustering [17]. The $SC$ assesses how well the clusters are separated, how close the clusters are, and whether they are stable for overlap clustering. Suppose we have $n$ samples in dataset $D$ that could be divided into $k$ clusters: $C_{1},...,C_{k}$. For any two data points $x,~{}x^{\prime}\in D$, the compactness of the cluster to which $x$ belongs can be defined as
@@
where $d(x,x^{\prime})$ represents the distance between $x$ and $x^{\prime}$, $|C_{i}|$ is the number of data points in cluster $i$. The degree of separation between $x$ and other clusters can be computed as
$b(x)=\mathop{\min}_{C_{j}:j\neq i}\bigg{\{}\frac{\sum_{x^{\prime}\in C_{j}}d(x,x^{\prime})}{|C_{j}|}\bigg{\}}.$ (2)
And the $SC$ of $x$ is
$SC(x)=\frac{b(x)-a(x)}{\max\{a(x),b(x)\}}.$ (3)
@@
The $SC$ reflects the clustering quality of data points, but it does not evaluate the topology structure of a Mapper output graph. Figure 2 shows an example of different Mapper graphs on the same dataset. Figure 2 (a) has a higher $SC$ but a poor topological feature. Conversely, Figure 2 (b) has a lower $SC$ but a good topological structure. Figure 2 (a) shows that a small overlap $p$ makes one of the circles disconnected, and however, more separable clusters lead to a high $SC$. Therefore, a metric that evaluates both clustering and topological structure is needed. Although, the topological information encoded in Mapper has been well studied theoretically [15]. There is no practical metric to evaluate the topology structure of Mapper graphs. Most existing work evaluates the topology structure of Mapper graphs based on the clustering results only. The extended persistence diagram has been proven to be a powerful tool for capturing topological signatures of Mapper graph [18]. In this manuscript, we alleviate it to provide a simple quantitative metric for evaluating the topology structure of Mapper output.
@@
Figure 2: An example of the classic Mapper algorithm on the same dataset but outputs different graphs. The dataset is shown in Figure 6 (a), the classic Mapper is implemented, and the clustering algorithm is DBSCAN with a radius of $0.5$ and a minimum of samples $3$. (a) The output graph when $n=12,p=0.01$ and its $SC=0.283$, $SC_{adj}=0.521$. This graph has a higher $SC$ but a poor topological structure. (b) The output graph when $n=12,p=0.1$ and its $SC=0.246$, $SC_{adj}=0.812$. This graph has a lower $SC$ but a good topological structure.
@@
The extended persistence diagram is an extension of the persistence diagram, the extended persistence diagram uses both sub-level sets and super-level sets as a filtration, therefore it contains more information than the original persistence diagram [19]. In practice, given a graph with a function defined on its nodes, we can compute the extended persistence diagram that reflects the topological features of this graph. For a Mapper graph, the function on a node can be naturally defined as the mean value of points in the node. The points on the diagram can be regarded as signatures of the Mapper graph. However, these signatures are not always meaningful, points near the diagonal are often seen as noise. To distinguish the real signal from the noise, one simple yet effective method is to calculate the confidence interval using bottleneck bootstrap and then identify the noise based on this interval [9]. The bottleneck bootstrap is an effective way to compute confidence set on a persistence diagram [20]. It uses bootstrap samples to get a bootstrap Mapper graph and then calculates the bottleneck distance from the original Mapper graph. Repeat this step many times, and an approximate distribution of bottleneck distances can be obtained, then we can easily calculate the confidence set with this approximate distribution. See Algorithm 4 in Appendix A for more details of the bottleneck bootstrap algorithm.
@@
We introduce a metric to evaluate the quality of the extended persistence diagram depending on the confidence set; we call this metric the topological signal rate, denoted as $TSR$. The $TSR$ is defined as the number of real signal points divided by the total number of points on the extended persistence diagram, serving as a quantitative indicator of the quality of the extended persistence diagram.
@@
An example of an extended persistence diagram and its $TSR$ is illustrated in Figure 3. The gray area is the confidence set estimated by the bottleneck bootstrap. Points inside are noises, and those outside are signal points. The $TSR$ in this example is $0.25$, indicating poor quality of the corresponding Mapper graph. To get a proper evaluation of the Mapper graph, we can combine the $TSR$ with $SC$ through weighted averaging. We call this metric an adjusted silhouette coefficient.
Figure 3: An example of the extended persistence diagram. There are a total of $8$ points in the diagram; the gray area is computed by the bottleneck bootstrap, and points inside this area are noise. Thus, $6$ points are noises, $2$ points are signals, and the $TSR$ is $0.25$.
@@
Figure 2 compares the value of $SC_{adj}$ with the original $SC$. The original $SC$ overlooks the topological structure of the Mapper graph, while the $SC_{adj}$ can give a reasonable evaluation. In the next section, we show more examples to validate the effectiveness of $SC_{adj}$.